OpenStudy (anonymous):

f(x)=x^7 -3, then f(x+h)

hartnn (hartnn):

just replace 'x' by 'x+h' and expand

OpenStudy (anonymous):

put x = x+h

hartnn (hartnn):

expand only if necessary, else leave your answer in the 7th exponent form

OpenStudy (anonymous):

i dont understand the expanding part

hartnn (hartnn):

don't expand. that will complicate it just replace x by x+h

OpenStudy (anonymous):

I did, not the right answer

hartnn (hartnn):

\(f(x+h)= (x+h)^7-3\) if your tool is showing this to be wrong, then either the question is incorrect or the tool is misbehaving

OpenStudy (uri):

f(x)=x^7 -3, then f(x+h) f(x)=(x+h)^7-3

hartnn (hartnn):

hope you got it right :)